# Sample, estimate, aggregate: A recipe for causal discovery foundation models

### v2

@@ 5.2 Low-cost adaptation
On real data from unknown data-generating processes, the ability to run causal discovery with different assumptions is useful when coupled with unsupervised methods for model selection (Faller et al., 2023). Table 3 illustrates two such approaches for continuous optimization algorithms. Methods like Avici, which operate on raw data, must re-train/finetune the model on the expanded data (around 4 days). In contrast, Sea can be run directly with a different estimation algorithm, or finetuned on the same data for around 4-6 hours (1 A6000) and $<4$ GB of GPU memory.
5.3 Limitations
Some empirical limitations of our implementations include: 1) a hard-coded maximum of 1000 variables (arbitrary), 2) generalize poorly to synthetic cyclic data, 3) err on the side of sparsity on real data, 4) inverse covariance can be numerically unstable on larger graphs (recommend models finetuned for correlation here), and 5) training requires full precision. These can be seen as artifacts of our training set and design choices.
@@ Acknowledgements
We thank Bowen Jing, Felix Faltings, Sean Murphy, and Wenxian Shi for helpful discussions regarding the writing; as well as Jiaqi Zhang, Romain Lopez, Caroline Uhler, and Stephen Bates for helpful feedback regarding the framing of this project.
@@
Appendix A Proofs and derivations
@@ A.2 Resolving marginal estimates into global graphs
{algorithm}
[…]
@@
Let $G=(V,E)$ be a directed acyclic graph with maximum degree $d$. For $S\subseteq V$, let $E^{\prime}_{S}$ denote the marginal estimate over $S$. Let $\mathcal{S}_{d}$ denote the superset that contains all subsets $S\subseteq V$ of size at most $d$. Algorithm A.4 maps $\{E^{\prime}_{S}\}_{S\in\mathcal{S}_{d+2}}$ to a pattern $E^{\prime}$ faithful to $G$.
@@
In the PC algorithm (Spirtes et al. (2001), A.1), its derivatives, and Algorithm A.4, there is no need to consider separating sets with cardinality greater than maximum degree $d$, since the maximum number of independence tests required to separate any node from the rest of the graph is equal to number of its parents plus its children (due to the Markov assumption).
@@
$\Rightarrow$ If $\{i,j\}\not\in C*$, then there must exist a separating set $S$ in $G$ of at most size $d$ such that $i\perp\!\!\!\perp j\mid S$. Then $S\cup\{i,j\}$ is a set of at most size $d+2$, where $\{i,j\}\not\in C^{\prime}_{S\cup\{i,j\}}$. Thus, $\{i,j\}$ would have been removed from $C^{\prime}$ in line 6 of Algorithm A.4.
@@
The PCI algorithm orients v-structures $i\rightarrow j\leftarrow k$ in $E^{*}$ if there is an edge between $\{i,j\}$ and $\{j,k\}$ but not $\{i,k\}$; and if $j$ was not in the conditioning set that removed $\{i,k\}$. Algorithm A.4 orients v-structures $i\rightarrow j\leftarrow k$ in $E^{\prime}$ if they are oriented as such in any $E^{\prime}_{S}$; and if $\{i,j\},\{j,k\}\in E^{\prime},\{i,k\}\not\in E^{\prime}$
@@
Given data that is faithful to $G$, Algorithm A.4 produces a pattern $E^{\prime}$ with the same connectivity and v-structures as $E^{*}$. Any additional orientations in both patterns are propagated using identical, deterministic procedures, so $E^{\prime}=E^{*}$. ∎
@@
A.3 Computational power of the axial attention model
@@
Existing literature on the universality and computational power of vanilla Transformers (Yun et al., 2019; Pérez et al., 2019) rely on generous assumptions regarding depth or precision. Here, we show that our axial attention-based model can implement the specific reasoning required to resolve marginal estimates under realistic conditions. In particular, we show that three blocks can recover the skeleton and v-structures in $O(N)$ width, and additional blocks have the capacity to propagate orientations. We first formalize the notion of a neural network architecture’s capacity to “implement” an algorithm. Then we prove Theorem 3.4 by construction.
@@
We construct a stack of $L\geq 3$ axial attention blocks that implement Algorithm A.4.
@@
Information theoretically, it should be possible to encode the same information in $\log N$ space, and achieve $O(\log N)$ width. For ease of construction, we have allowed for wider networks than optimal.
On the other hand, if we increase the width and encode each edge symmetrically, e.g. $(e_{i},e_{j},e_{j},e_{i}\mid i,j,j,i)$, we can reduce the number of blocks by half, since we no longer need to run each operation twice. However, attention weights scale quadratically, so we opted for an asymmetric construction.
@@ A.4 Robustness and stability
    We observe that global statistics can be estimated reliably in low data scenarios. For example, Figure 5 suggests that 200 examples suffice to provide a robust estimate over 100 variables in our synthetic settings. Therefore, even if the marginal estimates are erroneous, the neural network can learn the skeleton from the global statistics.
@@
    Most classical causal discovery algorithms are not stable with respect to edge orientation assignment. That is, an error in a single edge may propagate throughout the graph. Empirically, we observe that the majority vote of Gies achieves reasonable accuracy even without any training, while Fci suffers in this assessment (Table LABEL:table:synthetic-edge). However both Sea (Gies) and Sea (Fci) achieve high edge accuracy. Therefore, while the underlying algorithms may not be stable with respect to edge orientation, our pretrained aggregator seems to be robust.
@@ B.2 Baseline details
Deci (Geffner et al., 2022) was trained on all of the test datasets using their published hyperparameters. However, on all $N=100$ cases, the model failed to produce any meaningful results (adjacency matrices nearly all remained 0s with AUCs of 0.5). Thus, we only report results on $N=10,20$.
@@
Avici (Lorch et al., 2022) was run on all test datasets using the authors’ pretrained scm-v0 model, recommended for “arbitrary real-valued data.” Note that this model is different from the models described in their paper (denoted Avici-L and Avici-R), as it was trained on all of their synthetic data, including test sets. We sampled 1000 observations per dataset uniformly at random, with their respective interventions (the maximum number of synthetic samples used in their original paper), except for Sachs, which used the entire dataset (as in their paper). Though the authors provided separate weights for synthetic mRNA data, we were unable to use it since we did not preserve the raw gene counts in our simulated mRNA datasets.
@@
Corr and D-Corr are computed similarly, using global correlation and distance correlation, respectively (See C.2 for details).
B.3 Neural network design

### v3

@@ 5.2 SEA adapts to new data assumptions with zero to minimal finetuning
On real data from unknown distributions, these two strategies enable the ability to run causal discovery with different assumptions, which may be coupled with unsupervised methods for model selection (Faller et al., 2023). Table 3 illustrates this idea using the Sachs proteomics dataset. Sea can be run directly with a different estimation algorithm (FCI with polynomial kernel “Kci”), or finetuned for around 4-6 hours on 1 A6000 and $<4$ GB of GPU memory (correlation “Corr”). In contrast, methods like Avici must simulate new datasets based on each new assumption and re-train/finetune on these data (reportedly around 4 days).
5.3 SEA respects identifiability theory
While the identifiability of specific causal models is not a primary focus of this work, we show that Sea still respects classic identifiability theory. Specifically, while linear Gaussian models are known to be unidentifiable, Table 1 might suggest that both Sea and Dcdi perform quite well on these data – better than would be expected if graphs were only identifiable up to their Markov equivalence classes. This empirical “identifiability” may be the consequence of two findings. Common synthetic data generation schemes tend to result in marginal variances that reflect topological order (Reisach et al., 2021), and in additive noise models, it has been shown that marginal variances that are the “same or weakly monotone increasing in the [topological] ordering” result in uniquely identifiable graphs (Park, 2020). Data standardization can eliminate these artifacts of synthetic data generation. In Table 5.3, we see that after standardizing linear Gaussian data, our model performs no better than randomly selecting a graph from the Markov equivalence class (enumerated via pcalg (Kalisch et al., 2012)). The classic FCI algorithm is unable to reach this upper bound, suggesting that the amortized inference framework allows us to perform better in finite datasets.
5.4 Ablation studies
In addition to high performance and flexibility, one of the hallmarks of foundation models is their ability to act as few-shot learners when scaling up (Brown et al., 2020). We first confirm that Sea is indeed data-efficient, requiring only around 300-400 examples for performance to converge on datasets of $N=100$ variables, and outperforms inverse covariance (computed with 500 examples) at only 200 examples (Figure 4A). To probe for how this behavior emerges, we trained a “tiny” version of Sea (Gies) on approximately a quarter of the training data ($N=10,20$ datasets, 64.8 million examples). The tiny model performs nearly as well as the original on $N=10$ datasets when provided $T=50$ batches, but exhibits much poorer few-shot behavior with only $T=10$ batches (Figure 3). This demonstrates that Sea is able to ingest large amounts of data, leading to promising few-shot behavior.666 Due to computational limitations, we were unable to train larger models, as our existing training set requires several hundred GB in memory, and our file system does not support fast dynamic loading.
We also ablate each parameter of the estimation step to inform best practices. The trade-off between the number and size of batches may be relevant to estimation algorithms that scale poorly with the number of examples, e.g. kernel-based methods (Zhang et al., 2011). When given $T=100$ batches, Sea reaches reasonable performance at around 250-300 examples per batch (Figure 4B). Figure 4C further illustrates that on the harder Sigmoid datasets, 5 batches of size $b=500$ are roughly equivalent to 100 batches of size $b=300$. Finally, increasing the number of variables in each subset has minimal impact (Figure 4D), which is encouraging, as there is no need to incur the exponentially-scaling runtimes associated with larger subsets.
Finally, we analyze the impact of removing marginal estimates or global statistics (Table 8). First, we take a fully pretrained Sea (Gies) and set the corresponding hidden representations to 0. Performance drops more when $h^{\rho}$ is set to 0, indicating that our pretrained aggregator relies more on global statistics, though a sizable gap emerges in both situations. Then, we re-train Sea (Gies) on the $N=10$ datasets, with and without global statistics, so that lack of $\rho$ is in-distribution for the latter model, and the training sets are comparable. Here, we see that the “no $\rho$” version with $T=50$ estimates is on par with the original architecture with $T=10$ estimates, so the global statistic is equivalent to $\sim 40$ estimates. This roughly aligns with the theory that global statistics can expedite the skeleton discovery process (Section 3.4), as the number of estimates required to discover the skeleton of a $N=10$ graph is approximately (${10\choose 2}=45$ (Prop. A.8).
6 Discussion
Interventional experiments have formed the basis of scientific discovery throughout history, and in recent years, advances in the life sciences have led to datasets of unprecedented scale and resolution (Replogle et al., 2022; Nadig et al., 2024). The goal of these perturbation experiments is to extract causal relationships between biological entities, such as genes or proteins. However, the sheer size, sparsity, and noise level of these data pose significant challenges to existing causal discovery algorithms. Moreover, these real datasets do not fit cleanly into causal frameworks that are designed around fixed sets of data assumptions, either explicit (Spirtes et al., 1995) or implicit (Lorch et al., 2022). In this work, we approached these challenges through a causal discovery “foundation model.” Central to this concept were three goals. First, this model should generalize to unseen datasets whose data-generating mechanisms are unknown, and potentially out-of-distribution. Second, it should be easy to steer the model’s predictions with inductive biases about the data. Finally, scaling up the model should lead to data-efficiency. We proposed Sea, a framework that yields causal discovery foundation models. Sea was motivated by the idea that classical statistics and discovery algorithms provide powerful descriptors of data that are fast to compute and robust across datasets. Given these statistics, we trained a deep learning model to reproduce faithful causal graphs. Theoretically, we demonstrated that it is possible to infer causal graphs consistent with correct marginal estimates, and that our model is well-specified with respect to this task. Empirically, we implemented two proofs of concept of Sea that perform well across a variety of causal discovery tasks, easily incorporate inductive biases when they are available, and exhibit excellent few-shot behavior when scaled up.
While Sea provides a high-level framework for supervised causal discovery, there are several empirical limitations of the two implementations describe in this paper. These include: 1) an arbitrary, hard-coded maximum of 1000 variables, 2) poor generalization to synthetic cyclic data, 3) erring on the side of sparsity on real data, 4) numeric instability of inverse covariance on larger graphs, and as a result, 5) training requires full precision. The first three aspects may be addressed by modifying the architecture and/or synthetic training datasets, while the latter two can be addressed with more numerically stable statistics, like correlation.
More broadly, the success of supervised causal discovery algorithms derives from the fidelity of the data simulation procedure. In this paper, we present proofs of concept for the modeling framework, but we do not solve the problem of simulating realistic data. Of the data that may exist in the real world, the training data used here represent only a small, perhaps unrealistic, fraction. To achieve any semblance of trustworthiness on real applications, it is crucial to study the characteristics of each domain in detail – including common graph topologies and functional forms (Aguirre et al., 2024); the degree and nature of missing data (Hicks et al., 2018); and sources of measurement error or other covariates (Tran et al., 2020). When possible, we recommend that any insights be triangulated with other sources of knowledge. For example, while this work does not directly provide an inference framework, the predicted structure could be used to parametrize a generative model, whose likelihood could be evaluated on held-out interventions (Hägele et al., 2023). It is also important to check whether the inferred relationships compound upon any existing biases in the data. This is particularly important for sensitive domains like healthcare or legal applications.
This work also opens several directions for further investigation. The framework we describe utilizes a single causal discovery algorithm and a single global statistic. Classic causal discovery algorithms leverage diverse insights for identifying causal structure, e.g. the non-Gaussianity of noise (Shimizu et al., 2006) vs. conditional independence (Spirtes et al., 1995). Thus, different discovery algorithms or summary statistics may reveal different aspects of the causal structure. Learning to resolve these potentially conflicting views remains unexplored, both from experimental and theoretical perspectives. Furthermore, this work only shows that the axial attention architecture is well-specified as a model class and probes generalization empirically. This motivates theoretical studies into supervised causal discovery with regards to what information can be provably learned, e.g. in the style of of PAC (probably approximately correct) learning (Allen-Zhu et al., 2019).
In summary, we hope that this work will inspire a new avenue of research into causal discovery algorithms that are applicable to and informed by real applications.
@@ Acknowledgements
We thank Bowen Jing, Felix Faltings, Sean Murphy, and Wenxian Shi for helpful discussions regarding the writing; as well as Jiaqi Zhang, Romain Lopez, Caroline Uhler, and Stephen Bates for helpful feedback regarding the framing of this project. Finally, we thank our action editor Bryon Aragam and our anonymous reviewers for their invaluable suggestions towards improving this paper during the review process.
@@
Appendix A Theoretical motivations
@@ A.2 Resolving marginal estimates into global graphs
Algorithm 1 Resolve marginal estimates of $f\in\mathcal{F}$
1:  Input: Data $\mathcal{D}_{G}$ faithful to $G$
2:  Initialize $E^{\prime}\leftarrow K_{N}$ as the complete undirected graph on $N$ nodes.
3:  for $S\in\mathcal{S}_{d+2}$ do
4:     Compute $E^{\prime}_{S}=f(\mathcal{D}_{G[S]})$
5:     for $(i,j)\not\in E^{\prime}_{S}$ do
6:        Remove $(i,j)$ from $E^{\prime}$
7:     end for
8:  end for
9:  for $E^{\prime}_{S}\in\{E^{\prime}_{S}\}_{\mathcal{S}_{d+2}}$ do
10:     for v-structure $i\rightarrow j\leftarrow k$ in $E^{\prime}_{S}$ do
11:        if $\{i,j\},\{j,k\}\in E^{\prime}$ and $\{i,k\}\not\in E^{\prime}$ then
12:           Assign orientation $i\rightarrow j\leftarrow k$ in $E^{\prime}$
13:        end if
14:     end for
15:  end for
16:  Propagate orientations in $E^{\prime}$ (optional).
@@
Let $G=(V,E)$ be a directed acyclic graph with maximum degree $d$. For $S\subseteq V$, let $E^{\prime}_{S}$ denote the marginal estimate over $S$. Let $\mathcal{S}_{d}$ denote the superset that contains all subsets $S\subseteq V$ of size at most $d$. Algorithm 1 maps $\{E^{\prime}_{S}\}_{S\in\mathcal{S}_{d+2}}$ to a pattern $E^{\prime}$ faithful to $G$.
@@
In the PC algorithm (Spirtes et al. (2001), A.1), its derivatives, and Algorithm 1, there is no need to consider separating sets with cardinality greater than maximum degree $d$, since the maximum number of independence tests required to separate any node from the rest of the graph is equal to number of its parents plus its children (due to the Markov assumption).
@@
$\Rightarrow$ If $\{i,j\}\not\in C*$, then there must exist a separating set $S$ in $G$ of at most size $d$ such that $i\perp\!\!\!\perp j\mid S$. Then $S\cup\{i,j\}$ is a set of at most size $d+2$, where $\{i,j\}\not\in C^{\prime}_{S\cup\{i,j\}}$. Thus, $\{i,j\}$ would have been removed from $C^{\prime}$ in line 6 of Algorithm 1.
@@
V-structures are oriented $i\rightarrow j\leftarrow k$ in $E^{*}$ if there is an edge between $\{i,j\}$ and $\{j,k\}$ but not $\{i,k\}$; and if $j$ was not in the conditioning set that removed $\{i,k\}$. Algorithm 1 orients v-structures $i\rightarrow j\leftarrow k$ in $E^{\prime}$ if they are oriented as such in any $E^{\prime}_{S}$; and if $\{i,j\},\{j,k\}\in E^{\prime},\{i,k\}\not\in E^{\prime}$
@@
Given data that is faithful to $G$, Algorithm 1 produces a pattern $E^{\prime}$ with the same connectivity and v-structures as $E^{*}$. Any additional orientations in both patterns are propagated using identical, deterministic procedures, so $E^{\prime}=E^{*}$. ∎
@@
A.3 Model specification
Model specification and universality
In classical statistics, a statistical model can be expressed as a pair $(S,P_{\theta})$ for $\theta\in\Theta$, where $S$ is the sample space, $P$ is the family of distributions on $S$, and $\Theta$ is the space of parameters (McCullagh, 2002). Let $x_{1},\dots,x_{N}$ be observations of i.i.d. random variables in $S$, and let $P^{*}$ denote their common distribution. A statistical model is well-specified if $P^{*}=P_{\theta}$ for some $\theta\in\Theta$. Identifying the most appropriate class of models has long been an area of interest in statistics (Akaike, 1973). For simple models like linear regression, there are many ways to test for model specification, where the alternative may be that the data follow a non-linear relationship (Davidson & MacKinnon, 1981).
In machine learning, however, even the simplest architectures vary immensely. Whether a neural network is “well-specified” depends on many aspects, such as its width (number of hidden units), depth (number of layers), activation functions (sources of non-linearity), and more. Thus, instead of testing whether each neural network architecture is well-specified for each experiment, it is more common to show universality (or lack thereof). That is, given a class of functions $\mathcal{F}$ and the space of parameters $\Theta$, for every $f\in\mathcal{F}$ does there exist a $\theta\in\Theta$ that allows the neural network to approximate $f$ to arbitrary accuracy? For example, the most well-cited work for the universality of multi-layer perceptrons (Hornik et al., 1989) showed that:
“standard multilayer feedforward networks with as few as one hidden layer using arbitrary squashing functions are capable of approximating any Borel measurable function from one finite dimensional space to another to any desired degree of accuracy, provided sufficiently many hidden units are available.”
Note that this definition excludes neural networks that use rectified linear units (ReLU is unbounded), whose variants are ubiquitous today, as well as neural networks with bounded width. Universality under alternate assumptions has been addressed by many later works, e.g. Maiorov & Pinkus (1999); Shen et al. (2022).
With respect to modern Transformer architectures (Vaswani et al., 2017), the existing literature are similarly subject to constraints. For example, Pérez et al. (2019) shows that Transformers equipped with positional encodings are Turing complete, given infinite precision. Alternatively, Yun et al. (2020) proves that Transformers can approximate arbitrary continuous “sequence-to-sequence” functions, but requires in the worst case, exponential depth. Finally, Sanford et al. (2024) shows that one-dimensional self attention is unable to detect arbitrary triples along a sequence, unless the depth scales linearly as the input size (sequence length). This last case is particularly relevant to our setting, as v-structures are relations of three nodes, but even modern large language models “only” contain $\sim 100$ layers (e.g. GPT-4 is reported to have 120 layers). In general, these works present constructive proofs of either $\theta$ (Pérez et al., 2019; Yun et al., 2020) or pathological cases that cannot be handled (Sanford et al., 2024).
In this paper, we do not consider universality in a general sense, but rather limit our scope to the algorithms presented in Section A.2. This will allow us to make more realistic assumptions regarding the model class.
Causal identifiability for continuous discovery algorithms
Identifiability is central to causal discovery, as it is important to understand the degree to which a causal model can or cannot be inferred from data. In contrast to classic approaches, causal discovery algorithms that rely on continuous optimization must consider several additional aspects with respect to identifiability – namely, their objective; optimization dynamics; and model specification. We discuss representative works in light of these considerations.
First, the optima of the objectives must correspond to true graphs. NoTears (Zheng et al., 2018) focuses on the linear SEM case with least-squares loss, and they cite earlier literature regarding the identifiability of this setup (van de Geer & Bühlmann, 2013; Loh & Bühlmann, 2014; Aragam et al., 2015). Dcdi (Brouillard et al., 2020) proves that the graph that maximizes their proposed score is $\mathcal{I}$-Markov equivalent to the ground truth, subject to faithfulness and other regularity conditions. Since amortized causal discovery algorithms tend to be trained on data from a variety of data-generating processes, the identifiability of an arbitrary graph is less clear. Ke et al. (2023) cites Eberhardt et al. (2006) and claims that any graph is identifiable “in the limit of an infinite amount of [single-node hard] interventional data samples.” In our interventional setting, this also holds. Avici (Lorch et al., 2022) focuses solely on inferring graphs from data, without addressing whether they are identifiable. If we suppose that the graphs are identifiable, then in all three cases, the objectives can be written in terms of the KL divergence between the true and predicted edge distributions, which reaches its minimum when the predicted graph matches the true graph.
Second, the optimization process must not only converge to an optimum, but also (for amortized models) generalize to unseen data. This is, in general, difficult to show. Both NoTears and Dcdi acknowledge that due to the non-convexity of this optimization problem, the optimizer may converge to a stable point, which is not necessarily the global optimum. Avici includes an explicit acyclicity constraint, so they additionally cite Nandwani et al. (2019); Jin et al. (2020) for inspiration in constrained optimization of a neural network. Regarding generalization, existing results on what can be “provably” learned by neural networks are limited to simple architectures and/or algorithms (Allen-Zhu et al., 2019; Shao et al., 2022). Instead, both Avici and our work assess generalization empirically, by holding out certain causal mechanisms from training.
Finally, the implementation of each model must be well-specified. Since NoTears assumes linearity, it directly optimizes the (weighted) adjacency, and the only concern is whether linearity holds. Dcdi cites that deep sigmoidal flows are universal density approximators (Huang et al., 2018), which in turn invokes the classic result that a multilayer perceptron with sigmoid activations is a universal approximator (Cybenko, 1989). Avici does not discuss this aspect of their architecture, which is also based on sparse attention, but is otherwise quite different from ours. The following section proves by construction that our model is well-specified.
Axial attention architecture is well-specified
@@ A.3 Model specification
In the context of Section A.2, we show that three axial attention blocks (model depth) are sufficient to recover the skeleton and v-structures in $O(N)$ width, and we require $O(L)$ to propagate orientations along paths of length $L$. In the following section, we first formalize the notion of a neural network architecture’s capacity to “implement” an algorithm. Then we prove Theorem 3.4 by construction.
@@
That is, a single model implements an algorithm $f$ if for every input to $f$, the model outputs the corresponding output. A class of models parametrized by $\phi\in\Phi$ has the capacity to implement an algorithm if there exists at least one $\phi$ that implements $f$.
@@
We construct a stack of $L\geq 3$ axial attention blocks that implement Algorithm 1.
@@
Information theoretically, it should be possible to encode the same information in $\log N$ space, and achieve $O(\log N)$ width. For ease of construction, we have allowed for wider networks than optimal. On the other hand, if we increase the width and encode each edge symmetrically, e.g. $(e_{i},e_{j},e_{j},e_{i}\mid i,j,j,i)$, we can reduce the number of blocks by half, since we no longer need to run each operation twice. However, attention weights scale quadratically, so we opted for an asymmetric construction.
@@ A.4 Robustness and stability
    We observe that global statistics can be estimated reliably in low data scenarios. For example, Figure 4 suggests that 300 examples suffice to provide a robust estimate over 100 variables in our synthetic settings. Therefore, even if the marginal estimates are erroneous, the neural network can learn the skeleton from the global statistics.
@@
    Most classical causal discovery algorithms are not stable with respect to edge orientation assignment. That is, an error in a single edge may propagate throughout the graph. Empirically, we observe that the majority vote of Gies achieves reasonable accuracy even without any training, while Fci suffers in this assessment (Table 10). However both Sea (Gies) and Sea (Fci) achieve high edge accuracy. Therefore, while the underlying algorithms may not be stable with respect to edge orientation, our pretrained aggregator seems to be robust.
@@ B.2 Related work and baselines
AVICI (Lorch et al., 2022) is the most similar method to this work, though there are significant differences in both the causal discovery strategy and the implementation. Both works simulate diverse datasets for training, which differ in graph topology, causal mechanism, and type of exogenous noise; and both are attention-based architectures (Vaswani et al., 2017). The primary difference is that Avici operates over raw data, while we operate over summary statistics. While access to raw data may allow for richer modeling of relationships, it also increases the computational cost on large datasets. In fact, Avici’s model complexity scales quadratically as the number of samples and quadratically as the number of nodes. Our model does not explicitly depend on the number of data samples, but scales cubically as the number of nodes, as we attend over richer, pairwise features.
Both Avici and Sea are equivariant to the ordering of nodes. Avici is additionally invariant to the ordering of samples. Due to the two-track design of our aggregator, we also include message passing operations between the two input types (global statistics, marginal features), while Avici directly stacks Transformer layers. Finally, Avici explicitly regularizes for acyclicity. We follow the ENCO (Lippe et al., 2022) formulation (edges $i\to j$ and $j\to i$ cannot co-exist), since it is easier to optimize. Empirically, our predictions are still 99% acyclic (Table 18).
Avici was run on all test datasets using the authors’ pretrained scm-v0 model, recommended for “arbitrary real-valued data.” Note that this model is different from the models described in their paper (denoted Avici-L and Avici-R), as it was trained on all of their synthetic data, including test sets. We sampled 1000 observations per dataset uniformly at random, with their respective interventions (the maximum number of synthetic samples used in their original paper), except for Sachs, which used the entire dataset (as in their paper). Though the authors provided separate weights for synthetic mRNA data, we were unable to use it since we did not preserve the raw gene counts in our simulated mRNA datasets.
@@
DECI (Geffner et al., 2024) was trained on all of the test datasets using their published hyperparameters. However, on all $N=100$ cases, the model failed to produce any meaningful results (adjacency matrices nearly all remained 0s with AUCs of 0.5). Thus, we only report results on $N=10,20$.
@@
Corr and D-Corr are computed similarly, using global correlation and distance correlation, respectively (See C.1 for details).
B.3 Metrics
Throughout our evaluation, we compute metrics with respect to the ground truth graph. This means that in the observational setting, the “oracle” value of each metric will vary depending on the size of the equivalence class (e.g. if multiple graphs are observationally equivalent, the expected SHD is $>0$; see Section 5.3 for more analysis). While it is more correct to evaluate (observational) models with respect to the CPDAG that is implied by the predicted DAG, there were several reasons we also chose to evaluate with respect to the ground truth DAG.
  • This practice is quite common among continuous causal discovery algorithms, regardless of whether interventional data are available (Lorch et al., 2022; Brouillard et al., 2020; Hägele et al., 2023) or not (Sanchez et al., 2023). This may be primarily because their predicted graphs contain continuous probabilities, and metrics that reflect uncertainty (mAP, AUROC) are difficult to translate to the equivalence class.
  • It has been reported that common simulators produce “identifiable” linear Gaussian datasets, if the data are unstandardized (hence the VarSort baseline, Reisach et al. (2021)). This would mean that the MEC is actually smaller than expected. While it is possible to normalize all the data, it is unknown whether there are additional artifacts of the data generating process that may influence the (empirical) identifiability of the graphs, and thus, the “true” size of the equivalence class.
  • Our models and baselines are evaluated on the same graphs, but not all can incorporate interventional information. For comparability, we evaluated all predictions against the same ground truth, while noting that the equivalence class is larger in the observational case.
  • Fci, Gies, and VarSort don’t necessarily produce acyclic graphs, due to bootsrapping. In Appendix C, we also include the symmetric summary statistics as baselines, to quantify the “cost” of not directing edges, and of not using marginal estimates. In these cases, we cannot compute a corresponding CPDAG.
  • Practically, it was somewhat expensive to compute the CPDAG, especially for larger dense graphs ($N=100,E=400$ and beyond).
B.4 Training and hardware details